# Graph-Based Captioning: Enhancing Visual Descriptions by Interconnecting Region Captions

### v1

@@ Abstract
Humans describe complex scenes with compositionality, using simple text descriptions enriched with links and relationships. While vision-language research has aimed to develop models with compositional understanding capabilities, this is not reflected yet in existing datasets which, for the most part, still use plain text to describe images. In this work, we propose a new annotation strategy, graph-based captioning (GBC) that describes an image using a labelled graph structure, with nodes of various types. The nodes in GBC are created using, in a first stage, object detection and dense captioning tools nested recursively to uncover and describe entity nodes, further linked together in a second stage by highlighting, using new types of nodes, compositions and relations among entities. Since all GBC nodes hold plain text descriptions, GBC retains the flexibility found in natural language, but can also encode hierarchical information in its edges. We demonstrate that GBC can be produced automatically, using off-the-shelf multimodal LLMs and open-vocabulary detection models, by building a new dataset, GBC10M, gathering GBC annotations for about 10M images of the CC12M dataset. We use GBC10M to showcase the wealth of node captions uncovered by GBC, as measured with CLIP training. We show that using GBC nodes’ annotations—notably those stored in composition and relation nodes—results in significant performance boost on downstream models when compared to other dataset formats. To further explore the opportunities provided by GBC, we also propose a new attention mechanism that can leverage the entire GBC graph, with encouraging experimental results that show the extra benefits of incorporating the graph structure. Our datasets are released at https://huggingface.co/graph-based-captions.
Introduction
The availability of huge paired image/caption datasets has revolutionized our ability to produce joint vision-language embedding, paving the way for tasks like efficient caption-guided image generation within powerful multimodal foundation models [50, 36, 2, 38]. The quality and granularity of these datasets plays, therefore, a crucial role. While quality can be addressed by filtering out data [53, 17, 22] or, inversely, by improving caption quality through recaptioning [35, 14, 45, 16], there is ample interest in the community to provide more detailed, fine-grained information for each image [42, 3, 7, 15]. To obtain better annotations, we draw inspiration from compositionality, a fundamental characteristic of human perception that is reflected in the natural language used to describe our surroundings [12, 20, 30, 29, 4, 10]. Compositionality plays an especially important role when examining larger images found in the wild, which have a rich coarse-to-fine, hierarchical structure, commonly represented as a scene graph [31]. While scene graphs have been successfully applied to image retrieval [54, 31], generation [44, 18], and pre-training [26, 28], the scale of scene graph dataset is typically small. For instance, Visual Genome [33] only contains around k images.
@@ Introduction
1. We propose graph-based captioning (GBC), a new vision-language data format that captions images with a graph-based structure akin to scene graphs while retaining the flexibility and intuitiveness of plain text description. GBC contains four types of nodes: (1) an image node with captions of the entire image, (2) entity nodes that contain descriptions of individual objects, (3) composition nodes that link objects in the images of the same type, and (4) relation nodes that describe the spatial (“the tree is to the left of the tower”) or semantic (“The branch is covered in snow”) relationships between objects of different types (§ 3.1).
@@
2. We design a workflow to produce GBC annotations at scale. Inspired by recent recaptioning approaches [35, 14, 45, 16], we generate GBC annotations using the OSS LLaVA-1.6 [42, 43]. First, LLaVA generates short and long captions for the entire image, used to extract entities. Then, an object detection model (YOLO-World [9]) is employed to find bounding boxes for each entity. Subsequently, the same procedure is recursively invoked to produce a GBC for each proposal. Finally, LLaVA-1.6 is prompted to produce composition and relation captions that connect multiple entity nodes (§ 3.2).
@@
3. We create large-scale GBC dataset containing million images with words per image using that workflow. While ours is the first vision-language dataset that contains structured captions, a few recent datasets contain dense annotations, and only [65] has a scale that is similar to ours. Drawing on graph-encoders, we also design a baseline architecture to utilize this structure, using a new attention mechanism, SAHA (SAHA) (§ 4).
@@
4. We demonstrate experimentally in § 5 that the diversity of captions found in GBC nodes improves CLIP model performance across image-to-text retrieval, text-to-image retrieval, compositionality, and semantic segmentation tasks, while retaining comparative performance on zero-shot ImageNet classification. Our ablations highlight that it is not the density or scale of our dataset, but the structured nature itself, that leads to better performance. Remarkably, we observe that composition and relation nodes, which can only be obtained through the GBC workflow, boost performance. Finally, we perform ablation on the influence of annotation format on retrieval performance using a set-aside test set from GBC. In this case, we see that SAHA, when fed GBC annotations, provides comparable or even better performance than that obtained when describing an image with detailed captions, suggesting that GBC can be a promising alternative to traditional image captioning formats.
Related works
@@
In this section, we discuss related works on vision-language datasets. We refer the readers to Appendix A for works that are specific to CLIP [50] training.
@@
Improving image annotations with graph-based captioning
@@
3.1 Representing an image with graph-based captions
@@
An examplar GBC, generated automatically through our workflow is provided in Figure 1. It should be noted that the only manual addition in that graph comes from the "title label" of each node, which is obtained by taking the union of labels found in its incoming edges. Compared to the standard scene graph annotation, the use of node captions provides flexibility to describe complex concepts, while the underlying graph still captures the inherent structure of the image. Our dataset, whose construction is detailed in Section 3.2 next, includes several different types of captions tailored to the structure of the DAG. At the root image node, we provide both detailed and short captions to cater to varying levels of granularity. Captions at composition nodes and relation nodes explicitly describe the arrangement and interaction of multiple objects, while the captions at the entity nodes provide detailed description of a single object.
@@ 3.3 GBC1M and GBC10M
Following the process outlined in Section 3.2, we annotate the CC12M dataset [6] with graph-based captions using LLaVA-1.6 [42, 43] as the MLLM and Yolo-World [9] as the open-vocabulary detection model. Specifically, we construct two sets of annotations: GBC1M for a subset of around 1M of images, with all the queries performed with the Yi-34B version of LLaVA-1.6, and GBC10M for a subset of around 10M of images, with LLaVA-1.6 Yi-34B for image and composition queries, and LLaVA-1.6 Mistral-7B for entity and relation queries.222Our larger dataset does not cover the entire CC12M both because some images were no longer accessible at the time we accessed the images, and because we discard images for which the MLLM model’s reply to the image query does not comply with the prescribed format.
@@
We provide statistics of the above two datasets in Table 1. We note that these two datasets have very similar per-image statistics, with the number of words being the only exception, as LLaVA-1.6 Yi-34B tends to provide longer descriptions than LLaVA-1.6 Mistral-7B. Moreover, our datasets use an average number of around 500 words to describe each image. This is comparable to other dataset with rich annotations such as DCI (1111 words/img) [61] and DOCCI (136 words/img) [46]. We also compute the CLIP scores between the captions and their corresponding regions using the DFN-5B CLIP model [17], and we report their distribution for the GBC10M dataset in Figure 3. We note that the original CC12M caption achieves the highest CLIP scores, followed by the short synthetic caption for the entire image. This can be explained by the fact that in these two cases, the involved image-caption pairs more closely align with the training data of standard CLIP models.
Encoding GBC via structure-aware hierarchical attention
Alongside many ways to leverage GBC annotations, as we shall present in Section 5, we propose a simple text encoder architecture to incorporate structural information encoded in GBC graph along with node captions. Specifically, we present SAHA (SAHA) block which treats each caption as an individual sample, and introduces an additional cross-attention layer that enforces the captions to attend to their children.
Formally, we consider a caption graph with vertices and edges such that if and only if , , , and the label is included within the caption . In words, each vertex in the graph represents a caption from a node of the original graph and there is an edge from one caption to another only if the second caption describes part of the first caption. After tokenization of the captions, we can map the edge labels to a set of token positions of the source caption, which we write as . Then, the target caption annotates the source caption via the tokens at positions . Therefore, we can simply consider cross-attention with queries from these tokens and keys and values from the target caption. We illustrate this idea in Figure 4, where we zoom in on the additional cross-attention layer (SACA) on the right side of the figure. For completeness, we also provide the corresponding mathematical formula in Appendix D.
Our text encoder then stacks a number of SAHA blocks, effectively interleaving the vanilla self-attention layers that process, local, intra-caption information, with the SACA (SACA) layers that process global, inter-caption information in a structure-aware manner. Furthermore, the model acts as a classic text encoder in the absence of edges in the graph, i.e., when , dropping all SACA blocks.
Experiments
@@
We present in this section a comprehensive set of experiments to benchmark different image annotation schemes. Specifically, using CLIP model training as the main task, we show that GBC annotations can bring improvements on a range of benchmarks across classification, retrieval, and dense prediction tasks, compared to existing annotation schemes (Section 5.3). On retrieval tasks, we demonstrate how GBC allows one to encode denser, more descriptive textual information to better represent images as shown by the performance gain compared to existing annotation formats (Section 5.4). Missing experimental details are provided in Appendix E.
@@ 5.1 Annotation formats
Long caption. One can improve image description using a longer caption. We use long captions in our dataset. They are of 110 words on average, as compared to short captions, of only 28 words on average. We extend the context length of text encoders in CLIP models from 77 to 512 for this setup.
@@
Graph-based captions. Finally, we consider the GBC format as proposed in Section 3.1. The GBC format includes region captions, but also provides additional information, stored in relation and composition nodes. With this in mind, we explore three different ways to leverage GBC annotations:
@@
To fully benefit from the graph information, when available, we leverage the SAHA block proposed in Section 4. This allows us to encode the entire graph into text embeddings that also contain information from their respective subgraph. We refer to this method as GBC-graph.
As GBC annotation encapsulates all existing short, long, and region caption formats, we are able to instantiate all the setups by using only a subset of annotation available in our curated GMC10M dataset. Specifically, taking only the short or detailed caption at the root image node creates the short and long caption setup, respectively. To mimic the region caption setup, we drop the relation and composition captions from GBC annotations. By turning GBC into these settings, we ensure the same quality of text annotations across different methods.
@@ 5.2 Experimental setup
We perform CLIP training on our GBC10M dataset, while leaving out 10,151 samples as the test set. Following common practice, we use the CLIP score computed by a pre-trained CLIP model [17] to filter our training set, discarding the 5% of captions with the lowest scores for each type. In addition, we retain the original CC12M captions associated with each image. Specifically, in all setups, both the original caption and the short synthetic caption are consistently used as positive texts for the image during training. This prevents the severe distribution shifts that could occur from using only long or region captions when evaluating on standard benchmarks.
@@
Model and hyperparmeters. We use the standard CLIP ViT-B/16 model, with the only difference of longer context length of text encoder for long caption and GBC-concat, and a replacement of the vanilla transformer block by the SAHA block in text encoder for GBC-graph. We fix the global batch size (i.e., number of images in each batch) to 4,096 for all the methods. The models are trained for 45,000 steps with AdamW and cosine scheduler at a learning rate of . This roughly correspond to 20 epochs of training. We evaluate at the EMA checkpoint at epoch 10, as we observe that further training provides little to no improvement in performance across the benchmarks.
@@ 5.3 Evaluations on existing benchmarks
GBC annotation leads to clear performance gains by encoding relational information. Table 5.1 demonstrates that training models with more detailed textual information, such as long captions or region captions, consistently enhances downstream performance, particularly in retrieval tasks and dense prediction. However, the most significant improvements are seen with GBC-captions, which augment traditional region captions with relational and compositional descriptions. Given that the GBC workflow is uniquely positioned to provide these, this result demonstrates the soundness of GBC, capturing valuable insights not present in conventional captions.
@@ 5.4 Evaluation on GBC test set
To assess the effectiveness of different annotation formats, we provide the model with these annotations at test time. Annotations that better describe the images should, ideally, result in better retrieval performance when they are used. Specifically, we use our own test set and consider performing retrieval with the various types of annotations presented in Section 5.1. Note, however, that when using region captions or GBC-captions, no single text embedding can naturally encompass all the relevant information. To address this limitation, we perform retrieval based on the average CLIP score between the image embedding and the text embeddings of the provided captions in this setup. We report our results in Section 5.4, where the rows correspond to the annotations at training time and the columns correspond to the annotation at test time. Unsurprisingly, we see a strong tendency that when a model is trained with a certain annotation format, it performs the best when we use the same format for retrieval. Among the few exceptions, we note that models trained to pair with shorter captions may have better performance when concatenation of short captions is provided at test time. This leads us to the following two observations.
@@
Denser textual information improves retrieval performances. The table clearly shows that training with richer annotations—such as long captions, GBC-concat, or GBC-graph—enhances retrieval performance. This improvement suggests that these methods provide a more effective representation of the images. Specifically, GBC-graph yields the best performance, indicating that the proposed GBC format consists in a viable alternative to the commonly used detailed captions.

### v2

@@ Abstract
Humans describe complex scenes with compositionality, using simple text descriptions enriched with links and relationships. While vision-language research has aimed to develop models with compositional understanding capabilities, this is not reflected yet in existing datasets which, for the most part, still use plain text to describe images. In this work, we propose a new annotation strategy, graph-based captioning (GBC) that describes an image using a labeled graph structure, with nodes of various types. The nodes in GBC are created through a two-stage process: first, identifying and describing entity nodes; second, linking these nodes by highlighting compositions and relations among them. Since all GBC nodes hold plain text descriptions, GBC retains the flexibility found in natural language, but can also encode hierarchical information in its edges. We demonstrate that GBC can be produced automatically, using off-the-shelf multimodal LLMs and object detection models, by building a new dataset GBC10M that gathers GBC annotations for about 10M images of the CC12M dataset. Through CLIP training on GBC10M, we show that leveraging GBC nodes’ annotations—particularly those in composition and relation nodes—significantly boosts the model’s performance across various benchmarks compared to when other annotations are used. To further explore the opportunities provided by GBC, we also investigate the use of GBC as middleware for text-to-image generation, and show the extra benefits of incorporating the graph structure in this task.
1 Introduction
The availability of huge paired image/caption datasets has revolutionized our ability to develop advanced multimodal models, enabling a range of tasks such as efficient text-to-image synthesis, text-guided image manipulation, and fine-grained image understanding through multimodal large language models [68, 49, 2, 65, 56, 6]. The quality and granularity of these datasets plays, therefore, a crucial role. While quality can be addressed by filtering out data [71, 22, 30] or with a simple recaptioning strategy [48, 19, 60, 21], there is ample interest in the community to provide more detailed, fine-grained information for each image [56, 4, 9, 20]. To obtain better annotations, we draw inspiration from compositionality, a fundamental characteristic of human perception that is reflected in the natural language used to describe our surroundings [15, 27, 39, 38, 5, 13]. Compositionality plays an especially important role when examining larger images found in the wild, which have a rich coarse-to-fine, hierarchical structure, commonly represented by a scene graph [40]. While scene graphs have been successfully applied to image retrieval [72, 40], generation [59, 23], and pre-training [35, 37], the scale of scene graph dataset is typically small. For instance, Visual Genome [45] only contains around k images.
@@ 1 Introduction
1. We propose graph-based captioning (GBC), a new vision-language data format that captions images with a graph-based structure akin to scene graphs while retaining the flexibility and intuitiveness of plain text description. GBC contains four types of nodes: (1) an image node with captions of the entire image, (2) entity nodes that contain descriptions of individual objects, (3) composition nodes that link objects in the images of the same type, and (4) relation nodes that describe the spatial (“the tree is to the left of the tower”) or semantic (“The branch is covered in snow”) relationships between objects of different types. Importantly, encoding non-hierarchical relations as nodes rather than edges enables us to capture relationships involving more than two objects (§ 3.1).
@@
2. We design a workflow to produce GBC annotations at scale. Our approach combines a multimodal large language model (MLLM) with an open-vocabulary detection model. Initially, the MLLM generates both short and detailed captions for the entire image, which are used to identify entities. The detection model is then applied to locate bounding boxes for each identified entity. This process is recursively applied to create a GBC for each detected proposal. Finally, the MLLM is prompted to generate composition and relation captions that link multiple entity nodes together (§ 3.2).
@@
3. We create large-scale GBC dataset containing million images with words per image using the aforementioned workflow. While ours is the first vision-language dataset that contains structured captions, a few recent datasets contain dense annotations, and only [85] has a scale that is similar to ours. Our dataset is released under the CC BY-NC 4.0 license (§ 3.3).
@@
4. We demonstrate the benefit of GBC via CLIP training experiments. Concretely, we show that the diversity of captions found in GBC nodes improves CLIP model performance across image-to-text retrieval, text-to-image retrieval, compositionality, and semantic segmentation tasks, while retaining comparative performance on zero-shot ImageNet classification. Remarkably, we observe that composition and relation nodes, which can only be obtained through the GBC workflow, boost performance. Moreover, we perform ablation on the influence of annotation format on retrieval performance using a set-aside test set from GBC. In this case, we see that an architecture that is tailored to the GBC format provides comparable or even better performance than that obtained when describing an image with detailed captions, suggesting that GBC can be a promising alternative to traditional image captioning formats (§ 4).
5. We demonstrate the benefit of GBC as middleware for text-to-image generation. By breaking down the text-to-image generation process into the subtasks of text-to-GBC and GBC-to-image, we provide users with a powerful intermediary for fine-grained image manipulation. The text-to-GBC subtask is handled by a lightweight language model, while the GBC-to-image subtask can be completed through a training-free approach. Importantly, we show that the inclusion of graph information significantly enhances the performance of a baseline method that otherwise performs poorly when only bounding box information is used (§ 5).
2 Related works
@@
In this section, we discuss related works on vision-language datasets. We refer the readers to Appendix A for works that are specific to CLIP [68] and text-to-image models.
@@
3 Improving image annotations with GBC
@@
3.1 Representing an image with GBC
@@
An examplar GBC, generated automatically through our workflow is provided in Figure 1. It should be noted that the only manual addition in that graph comes from the "title label" of each node, which is obtained by taking the union of labels found in its incoming edges. Compared to the standard scene graph annotation, the use of node captions provides flexibility to describe complex concepts, while the underlying graph still captures the inherent structure of the image. Our dataset, whose construction is detailed in Sec. 3.2 next, includes several different types of captions tailored to the structure of the DAG. At the root image node, we provide both detailed and short captions to cater to varying levels of granularity. Captions at composition nodes and relation nodes explicitly describe the arrangement and interaction of multiple objects, while the captions at the entity nodes provide detailed description of a single object.
@@ 3.3 GBC1M and GBC10M
Following the process outlined in Sec. 3.2, we annotate the CC12M dataset [8] with graph-based captions using LLaVA-1.6 [56, 57] as the MLLM and Yolo-World [12] as the open-vocabulary detection model. Specifically, we construct two sets of annotations: GBC1M for a subset of around 1M of images, with all the queries performed with the Yi-34B version of LLaVA-1.6, and GBC10M for a subset of around 10M of images, with LLaVA-1.6 Yi-34B for image and composition queries, and LLaVA-1.6 Mistral-7B for entity and relation queries.222Our larger dataset does not cover the entire CC12M both because some images were no longer accessible at the time we accessed the images, and because we discard images for which the MLLM model’s reply to the image query does not comply with the prescribed format.
@@
We provide statistics of the above two datasets in Tab. 1. We note that these two datasets have very similar per-image statistics, with the number of words being the only exception, as LLaVA-1.6 Yi-34B tends to provide longer descriptions than LLaVA-1.6 Mistral-7B. Moreover, our datasets use an average number of around 500 words to describe each image. This is comparable to other dataset with rich annotations such as DCI (1111 words/img) [80] and DOCCI (136 words/img) [62].
4 CLIP training with GBC
@@
We present in this section a comprehensive set of experiments to compare different image annotation schemes from a CLIP training perspective. We first show that compared to existing annotation schemes, GBC annotations can bring improvements on a range of benchmarks across classification, retrieval, and dense prediction task. Then, we demonstrate how GBC allows one to encode denser, more descriptive textual information to better represent images on retrieval tasks. Additional experimental details and further ablations are respectively provided in Appendices E and F.
@@ 4.1 Annotation formats
Long caption. One can improve image description using a longer caption. The long captions that we use in our experiments are of 110 words on average, as compared to short captions, of only 28 words on average. We extend the context length of text encoders in CLIP models from 77 to 512 for this setup.
@@
Graph-based captions. Finally, we consider the GBC format as proposed in Sec. 3.1. The GBC format includes region captions, but also provides additional information, stored in relation and composition nodes. With this in mind, we explore three different ways to leverage GBC annotations:
@@
To fully benefit from the graph information, when available, we introduce additional cross-attention layers to leverage the graph topology (see Sec. D.1 for details). This allows us to encode the entire graph into a text embedding that also contains the information about the graph structure. We refer to this method as GBC-graph.
As GBC annotation encapsulates all existing short, long, and region caption formats, we are able to instantiate all the setups by using only a subset of annotation available in our curated GBC10M dataset. Specifically, taking only the short or detailed caption at the root image node creates the short and long caption setup, respectively. To mimic the region caption setup, we drop the relation and composition captions from GBC annotations. Additionally, to isolate the impact of using long captions and the additional captions uncovered by GBC, we omit long captions from both the region caption setup and the three GBC setups. By adapting GBC into these configurations, we ensure consistent text annotation quality across different methods.
@@ 4.2 Experimental setup
We perform CLIP training on our GBC10M dataset, while leaving out 10,151 samples as the test set. Following common practice, we use the CLIP score computed by a pre-trained CLIP model [22] to filter our training set, discarding the 5% of captions with the lowest scores for each type. In addition, we retain the original CC12M captions associated with each image. Specifically, in all setups, both the original caption and the short synthetic caption are used as positive texts for the image during training. This prevents the severe distribution shifts that could occur from using only long or region captions when evaluating on standard benchmarks.
@@
Model and hyperparmeters. We use the standard CLIP ViT-B/16 model, with the only difference of longer context length of text encoder for long caption and GBC-concat, and a replacement of the vanilla transformer block by our dedicated attention block in text encoder for GBC-graph. We fix the global batch size (i.e. number of images in each batch) to 4,096 for all the methods. The models are trained for 45,000 steps with AdamW and cosine scheduler at a learning rate of . This roughly correspond to 20 epochs of training. We evaluate at the EMA checkpoint at epoch 10, as we observe that further training provides little to no improvement in performance across the benchmarks.
@@ 4.3 Evaluations on existing benchmarks
GBC annotation leads to clear performance gains by encoding relational information. Table 4.1 demonstrates that training models with more detailed textual information, such as long captions or region captions, consistently enhances downstream performance, particularly in retrieval tasks and dense prediction. However, the most significant improvements are seen with GBC-captions, which augment traditional region captions with relational and compositional descriptions, yielding about a recall increase for retrieval tasks and a boost in mIOU in segmentation task compared to using only short captions. Given that the GBC workflow is uniquely positioned to provide such relational information, this demonstrates that GBC captures valuable insights not present in conventional captions.
@@ 4.4 Evaluation on GBC test set
To assess the effectiveness of different annotation formats, we provide the model with these annotations at test time. Annotations that better describe the images should, ideally, result in better retrieval performance when they are used. Specifically, we use our own test set and consider performing retrieval with the various types of annotations presented in Sec. 4.1. Note, however, that when using region captions or GBC-captions, no single text embedding can naturally encompass all the relevant information. To address this limitation, we perform retrieval based on the average CLIP score between the image embedding and the text embeddings of the provided captions in this setup. We report our results in Sec. 4.4, where the rows correspond to the annotations at training time and the columns correspond to the annotation at test time. Unsurprisingly, we see a strong tendency that when a model is trained with a certain annotation format, it performs the best when we use the same format for retrieval. Among the few exceptions, we note that models trained to pair with shorter captions may have better performance when concatenation of short captions is provided at test time. This leads us to the following two observations.
@@
Denser textual information benefits retrieval. The table clearly shows that training with richer annotations—such as long captions, GBC-concat, or GBC-graph—enhances retrieval performance. This improvement suggests that these methods provide a more effective representation of the images. Specifically, GBC-graph yields the best performance, indicating that the proposed GBC format consists in a viable alternative to the commonly used detailed captions.